### v3

# Adaptive Focus Memory for Language Models

Abstract.
Large language models (LLMs) are increasingly deployed in multi-turn dialogue settings, yet their behavior remains bottlenecked by naive history-management strategies. Replaying the full conversation at every turn is simple but costly, while recency-based truncation or static summarization often causes early, high-impact user constraints to drift out of effective context. As a result, models may retain text without reliably applying it when it matters.
We present Adaptive Focus Memory (AFM), a lightweight context management system that dynamically assigns each past message one of three fidelity levels—Full, Compressed, or Placeholder—based on semantic relevance, temporal decay, and importance classification. AFM packs messages chronologically under a fixed token budget, preserving critical constraints at high fidelity while allowing low-importance context to degrade gracefully.
We evaluate AFM on two multi-turn dialogue benchmarks designed to stress long-horizon constraint preservation: a safety-critical travel scenario involving a user with a severe peanut allergy, and a policy-critical tax compliance scenario involving an illegal evasion request. Under strict grading that requires both explicit constraint recall and appropriately conditioned generation, AFM succeeds in 83.3 percent of allergy runs where all baseline strategies fail under strict grading, and preserves correct refusal behavior on the tax benchmark.
These results demonstrate that effective dialogue memory requires more than retaining prior text. Selectively allocating fidelity across past messages enables reliable constraint preservation under bounded context growth, without modifying model weights or introducing external retrieval infrastructure. We release an open-source implementation of AFM compatible with OpenAI-style chat APIs to support reproducible research and practical deployment.
1. Introduction
Large language models (LLMs) such as ChatGPT, Claude, and Gemini have become powerful general-purpose dialogue agents. However, their performance remains constrained by limited context windows and the high inference-time cost of each token. As conversations grow in length, naïvely replaying the full history at every turn leads to bloated prompts, rising latency, and mounting API expenses—posing serious challenges for multi-agent systems, assistants, and long-form interactions.
To mitigate these inefficiencies, prior work has explored memory-management strategies like Retrieval-Augmented Generation (RAG) and summarization. RAG stores prior content in external vector databases and retrieves relevant chunks per query. Summarization periodically compresses old messages into short abstractions. However, both approaches face limitations in dialogue settings: RAG pipelines require dedicated retrieval infrastructure and often disrupt conversational flow, while static summaries introduce irreversible information loss and cannot adapt to shifting user intent.
Adaptive Focus Memory (AFM) introduces a dynamic alternative. Rather than replaying all history or summarizing indiscriminately, AFM allocates each past message to one of three fidelity levels: FULL (included verbatim), COMPRESSED (summarized via an LLM or heuristic), or PLACEHOLDER (replaced with a short reference stub). This selective replay preserves salient information while minimizing prompt length.
AFM scores messages using a combination of embedding-based semantic similarity, half-life recency decay, and an importance classifier. It can operate in fully offline (heuristic) or LLM-assisted modes, and integrates seamlessly into OpenAI-compatible chat pipelines. In benchmark evaluations on a safety-critical synthetic dialogue scenario, AFM reduces token usage by roughly two-thirds while preserving factual continuity.
Contributions. This work makes the following contributions:
  • We propose a dynamic context selection framework that scores past messages using semantic similarity, temporal recency, and LLM-based importance classification.
  • We introduce a multi-fidelity replay mechanism that assigns each message to FULL, COMPRESSED, or PLACEHOLDER form and packs messages chronologically under a fixed token budget.
  • We demonstrate that AFM preserves safety-critical constraints more reliably than naive replay and recency-based heuristics under strict evaluation.
  • We release an open-source AFM implementationhttps://github.com/cruz209/AFMforLLM suitable for reproducible research and integration into production systems.
The rest of the paper is organized as follows: Section 2 reviews prior work. Section 3 (Method) describes the AFM algorithm. Section 4 (Experiment) presents benchmarks. Section 5 (Discussion) covers limitations and broader implications, and Section 6 concludes.
2. Related Work
2.1. Long-Context Language Models
Early transformer architectures [2] established the self-attention mechanism that enables contextual reasoning but introduced a quadratic memory cost that continues to limit practical context length. Subsequent large-scale generative models such as GPT-3/4 [3, 4] and LLaMA [5] extend context windows but still pay linear inference-time cost proportional to prompt length. Recent industrial systems (e.g., Claude 3 [6]) expand length further, yet running them efficiently remains a bottleneck for multi-turn interaction.
Several works propose architectures designed for local adaptation or long-context specialization. introduce locally trainable memory layers for more efficient fine-tuning, while Geva et al. [11] show that transformer feed-forward layers operate implicitly as key–value memory stores. These approaches improve contextual capacity but do not directly reduce inference cost, which remains proportional to the entire visible context.
2.2. Retrieval-Augmented and External-Memory Approaches
Retrieval-Augmented Generation (RAG) [7] offloads knowledge into external vector stores and selectively injects relevant snippets at inference time. While effective for fact-intensive tasks, RAG often underperforms in conversational settings due to retrieval noise and lack of discourse continuity. Roberts et al. [10] demonstrate the limits of relying purely on internal parameters for knowledge storage, motivating hybrid external-memory architectures.
More recent surveys [8] highlight the difficulty of maintaining coherence when summarizing or retrieving historical dialogue, particularly when the user’s intent evolves across turns. Structured-memory architectures attempt to improve retrieval fidelity, yet these systems require substantial infrastructure—vector databases, indexing, and latency-heavy retrieval pipelines.
2.3. Summarization and Compression of Conversational History
A classical strategy for context control is periodic summarization, either extractive or abstractive. Wu et al. [9] examine long-context summarization to reduce token load, but fixed-frequency summarization introduces irreversible information loss and can cause hallucinations when summaries drift from user intent.
Extractive heuristics and sentence-selection methods remain lightweight but often fail to preserve subtle constraints—especially user-specific preferences or safety-critical details. Abstractive models perform better but at the cost of additional inference steps, which increases cumulative compute.
2.4. Tokenization and Efficient Inference
Inference cost is known to scale roughly linearly with input token count across major LLM stacks. Industry-standard tokenizers such as tiktoken [12] and dataset libraries such as HuggingFace Datasets provide the measurement tools used to benchmark compression efficacy in this work.
Compute-efficient LLM design groups (DeepSeek, OpenAI, Anthropic, Meta) increasingly emphasize token-level cost reduction as a primary engineering constraint. However, most existing work focuses on model-level changes (attention variants, KV caching, or speculative decoding), whereas AFM operates orthogonally: it reduces cost via dynamic, semantic-aware prompt packing without any model modifications.
2.5. Position of AFM Within the Literature
AFM differs from prior approaches in three key ways:
  • Dynamic fidelity assignment: Previous systems often treat all messages uniformly (full retention or global summary), whereas AFM assigns each message one of three fidelity levels—FULL, COMPRESSED, or PLACEHOLDER—based on semantic relevance, recency, and importance.
  • Model-agnostic, plug-and-play design: AFM requires no architectural changes, no retrieval store, and no fine-tuning. It operates entirely at the prompt layer and is implemented as a simple Python library.
Taken together, AFM occupies a space between RAG, KV-extended transformers, and classical summarization—providing a memory framework whose primary objective is token-efficient factual continuity rather than scaling raw context length or storing external facts.
3. Method
3.1. Overview.
Adaptive Focus Memory (AFM) is implemented as a pluggable context manager that operates entirely at the prompt-construction layer. Rather than replaying the full conversation history at every turn, AFM decides, for each past message, whether to include it verbatim, include a compressed summary, or replace it with a short placeholder stub.
We frame the problem as a heuristic, greedy packing process under a fixed budget . Given a conversation history $H=\{m_{1},\dots,m_{t}\}$, a current query $q_{t}$, and a maximum prompt budget $B$ (in tokens), AFM assigns each message $m_{i}$ a fidelity tier
$f(m_{i})\in\{\textsc{Full},\textsc{Compressed},\textsc{Placeholder}\}$
and then constructs a chronological prompt whose total estimated token length does not exceed $B$, whenever possible. The implementation prioritizes messages that are semantically similar to $q_{t}$, temporally recent, and classified as important by a small LLM-based classifier when an API is available.
3.2. Relevance scoring.
The reference implementation combines three signals in a single scalar score per message: semantic similarity, recency, and an LLM-based importance label.
Semantic similarity.
Each message $m_{i}$ is embedded once (lazily on first use) and cached. Given the embedding $E(m_{i})$ and the embedding $E(q_{t})$ of the current query, AFM computes cosine similarity
$\operatorname{sim}(m_{i},q_{t})=\frac{E(m_{i})\cdot E(q_{t})}{\|E(m_{i})\|\,\|E(q_{t})\|}.$
The reference implementation uses an OpenAI embedding model (text-embedding-3-small) when an API key is configured, and a dependency-free hashing-based embedder otherwise.
Recency weighting.
Let k=tik=t-i be the number of turns since message $m_{i}$ was created and let $h$ be a configurable half-life parameter. AFM computes a recency weight
$w_{\text{recency}}(m_{i})=0.5^{k/h},$
so that the influence of a message decays exponentially as it moves further back in the dialogue. In our experiments we use a fixed half-life $h=12$ turns unless otherwise noted.
Importance classification.
When an OpenAI client is available, AFM calls a small chat model (gpt-4o-mini) once per message to classify its importance as CRITICAL, RELEVANT, or TRIVIAL. The classifier is prompted to treat safety-sensitive facts (e.g., medical conditions) as CRITICAL. If no client is configured, all messages default to TRIVIAL and no classification calls are made. AFM does not evaluate the accuracy of the importance classifier, its behavior is prompt-driven.
Final scoring rule.
The actual scoring function in the reference implementation is piecewise, not a single multiplicative expression. Let $\operatorname{sim}_{i}$ be the cosine similarity and $w_{\text{recency},i}$ the recency weight for $m_{i}$. Then:
[…]
Critical messages are thus force-elevated to the maximum score, while relevant and trivial messages use similarity- and recency-weighted scores with different scales. These scores are only used for fidelity decisions; they do not reorder the conversation.
3.3. Fidelity assignment and context packing.
AFM uses the scores $s_{i}$ to assign an intended fidelity tier to each message before packing:
  • If $s_{i}\geq\tau_{\text{high}}$, mark $m_{i}$ as intended Full.
  • Else if $s_{i}\geq\tau_{\text{mid}}$, mark $m_{i}$ as intended Compressed.
  • Else, mark $m_{i}$ as intended Placeholder.
The thresholds $\tau_{\text{high}}$ and $\tau_{\text{mid}}$ are hyperparameters (e.g., $\tau_{\text{high}}=0.45$, $\tau_{\text{mid}}=0.25$ in our OpenAI-backed experiments).
Packing is then performed in strict chronological order. Let $B$ be the token budget and $b_{\text{left}}$ the remaining tokens. For each message $m_{i}$ (from oldest to newest), AFM attempts to include the highest-fidelity representation consistent with its intended tier and the remaining budget:
  • If $m_{i}$ is intended Full, AFM:
    1. (1)
      tries the full text; if the estimated token count fits within $b_{\text{left}}$, it is appended;
    2. (2)
      otherwise, generates or reuses a compressed summary and tries that;
    3. (3)
      if the summary still does not fit, it falls back to a stub.
  • If $m_{i}$ is intended Compressed, AFM attempts the compressed form first and falls back to a stub if the compressed form does not fit.
  • If $m_{i}$ is intended Placeholder, AFM only attempts the stub.
These hyperparameters were chosen heuristically based on a small number of
development conversations; no large-scale tuning was performed.
In all cases, AFM estimates token usage with a shared TokenCounter before committing to adding a representation. If even the stub for a message would exceed the remaining budget, that message is dropped from the prompt. Therefore, AFM attempts to maintain a cheap trace of low-importance turns, but it does not guarantee that every turn is represented under extremely tight budgets.
3.4. Compression layer.
The reference implementation supports two interchangeable compressors via a common Compressor interface:
  • HeuristicCompressor is a fully local, extractive compressor. It operates by ranking sentences using simple lexical overlap with a query hint, mild length penalties, and positional bias. It then truncates to a target token budget based on the shared TokenCounter. This compressor enforces the budget deterministically.
  • LLMCompressor is an abstractive compressor that calls an OpenAI chat model (gpt-4o-mini in our experiments) with a prompt instructing it to rewrite the input under a specified token budget. The compressor relies on the model to follow the budget instruction; the current implementation does not re-check or truncate the returned summary.
AFM selects the compressor at initialization time. In the reference code, OpenAI-backed runs use LLMCompressor; offline runs fall back to HeuristicCompressor.
3.5. Token budgeting and enforcement.
Token counts are estimated by a TokenCounter utility. When the tiktoken library is available, it uses the encoding for gpt-4o-mini and falls back to cl100k_base on error; if tiktoken is not available, it uses a naive whitespace-based word count as a proxy for token length.
During packing, AFM maintains a running budget $b_{\text{left}}$, initialized to the user-provided budget $B$. Every time it proposes to add a representation (full text, compressed summary, or stub), it first estimates the token count with TokenCounter. If the representation fits ($\text{tokens}\leq b_{\text{left}}$), AFM appends it and decrements $b_{\text{left}}$; otherwise, it tries a lower-fidelity form or drops the message entirely if no representation can fit.
A system preamble (if provided) is treated like any other message: it is included first, subject to the same budget accounting.
3.6. System architecture and API
AFM is implemented as a modular Python package centered on a FocusManager class. The key methods are:
  • add_message(role, content): append a new message to the internal history. Messages are assigned unique IDs, and embeddings are computed lazily.
  • build_context(current_query, budget_tokens, system_preamble): compute relevance scores, assign intended fidelity tiers, and return a list of $(\text{role},\text{content})$ pairs representing the packed prompt, along with summary statistics (e.g., total tokens used, number of compressed messages, etc.).
The same interface works in both online and offline modes and is designed to drop into OpenAI-style chat pipelines with minimal glue code.
3.7. Backend integration.
The reference implementation supports two embedding backends and two compression backends:
  • When an OpenAI API key is available, AFM uses an OpenAIEmbedder that wraps text-embedding-3-small and the LLMCompressor for abstractive compression.
  • When no API key is available, AFM falls back to a local HashingEmbedder and the HeuristicCompressor.
Similarly, importance classification is only performed when an OpenAI client is configured. If the classifier or compressor calls fail due to network or API errors, the current implementation allows the exception to propagate rather than attempting automatic retries or alternative backends.
3.8. Fallback and robustness.
AFM’s robustness features are deliberately minimal in the reference code. The only automatic “fallback” inside the core algorithm is fidelity downgrading: if a message cannot fit in Full form, AFM tries Compressed, then Placeholder. Beyond that, there is no automatic retry logic, rate limit handling, or circuit-breaking. Adding such mechanisms is left for future versions.
3.9. Design philosophy.
AFM is intentionally model-agnostic and infrastructure-light. It requires no vector database, no KV-cache modifications, and no fine-tuning. All state remains in process as a simple list of messages plus per-message metadata (embedding, score, compression state).
In contrast to static summarization approaches that periodically compress the entire history into a single buffer, AFM scores and revisits every message at each call to build_context. This allows it to adapt to changes in user intent over time while still providing explicit, budget-aware control over how much space each past turn occupies in the prompt.
4. Experiments
4.1. Objective and scope.
Our experiments evaluate whether AFM can reliably preserve high-importance constraints across multi-turn dialogue while controlling context growth. We focus on two safety-critical constraint types that appear frequently in production assistants but are not well covered by standard LLM benchmarks:
  1. (1)
    Personal safety constraints that must be remembered and applied later (e.g., a life-threatening food allergy).
  2. (2)
    Policy/legal constraints that must remain binding even after long benign context (e.g., refusing illegal tax-evasion assistance).
Both settings target a relatively new problem space: context management as a first-class system component. Most existing evaluations either (i) assume the full conversation history is always available, or (ii) test the model in isolation without stressing the memory policy that a deployed system actually uses. In contrast, we evaluate AFM as a history-management strategy under long, mixed-topic conversations where the binding constraint appears early and is not restated verbatim.
4.2. Why a custom benchmark.
We found that off-the-shelf safety or instruction-following benchmarks do not cleanly isolate the problem AFM addresses: constraint drift under realistic, long dialogue with imperfect memory. In particular, we need:
  • Constraint non-repetition: the constraint is stated once early, then later questions implicitly require it.
  • Mixed benign distractors: intervening turns that are plausible in real usage but semantically distant.
  • Strict pass semantics: it is not enough to merely mention the constraint; the response must treat it with appropriate seriousness.
Accordingly, we use two scripted, multi-turn scenarios with automatic grading and full transcript logging.
4.3. Task design.
We consider two benchmarks.
Peanut allergy travel benchmark.
A user planning a trip to Thailand states early: they have a severe peanut allergy. The conversation then proceeds through multiple benign topics (destinations, transport, cultural experiences), and culminates in a food recommendation query. A run passes if the final recommendation (i) recalls the allergy and (ii) conditions food guidance appropriately (e.g., avoiding peanut-heavy dishes, providing high-salience precautions).
Tax compliance benchmark.
A user provides an explicit constraint early: advice must comply with U.S. tax law and must not assist illegal evasion. After many benign tax questions, the user escalates to an explicit illegal request (hiding unreported freelance income). A run passes if the assistant clearly refuses and redirects toward legal, compliance-oriented alternatives.
4.4. Compared methods.
We compare AFM against three baselines corresponding to our benchmark scripts:
  1. (1)
    Default (stateless) chat. Only the current user turn (plus a fixed system prompt) is sent.
  2. (2)
    Naive truncated replay. Append prior turns until a global budget is hit, then drop oldest turns.
  3. (3)
    Recency-based compression. Keep the most recent turns verbatim; compress older turns to a fixed local budget.
AFM uses selective fidelity allocation via FocusManager: high-importance items are retained at Full fidelity, mid-importance items are compressed, and low-importance items are reduced to short stubs.
4.5. Implementation details.
All experiments use OpenAI’s gpt-4o-mini for generation. When enabled, the same model family is used for compression and importance classification. Token usage is measured with the shared tokenizer-backed TokenCounter. Latency is measured as wall-clock time from request submission to completion receipt.
4.6. Evaluation metrics.
Pass/fail semantics.
For both benchmarks, we adopt a strict binary criterion:
A run is a considered pass if and only if the model both (i) explicitly recalls the binding constraint and (ii) generates a response that treats it with appropriate seriousness.
This intentionally separates memory failures (constraint not recalled) from generation failures (constraint recalled but handled weakly, e.g., unsafe dish suggestions with soft disclaimers).
Reported metrics.
For each method and benchmark (30 seeds each), we report:
  • Pass rate (%).
  • Token usage on the graded turn (mean $\pm$ std).
  • Latency on the graded turn (mean $\pm$ std).
4.7. Main results.
Table 1 reports results on both benchmarks.
Method Pass rate Tokens (mean $\pm$ std) Latency (s, mean $\pm$ std)
Peanut allergy benchmark (Thailand travel)
Default (stateless) 0/30 (0%) 194.0 $\pm$ 0.0 4.04 $\pm$ 0.77
Naive truncated replay 0/30 (0%) 34.3 $\pm$ 46.5 11.49 $\pm$ 2.35
Recency compression 0/30 (0%) 33.9 $\pm$ 45.5 11.33 $\pm$ 1.90
AFM 25/30 (83.3%) 286.0 $\pm$ 0.0 21.24 $\pm$ 2.01
Tax compliance benchmark (illegal evasion request)
Default (stateless) 30/30 (100%) 44.0 $\pm$ 0.0 0.69 $\pm$ 0.17
Naive truncated replay 30/30 (100%) 32.0 $\pm$ 0.0 0.78 $\pm$ 0.21
Recency compression 30/30 (100%) 32.0 $\pm$ 0.0 0.75 $\pm$ 0.20
AFM 30/30 (100%) 286.0 $\pm$ 0.0 20.84 $\pm$ 2.61
Table 1. Live benchmark results (30 seeds per benchmark). A pass requires both constraint recall and appropriate constraint-conditioned generation. AFM substantially improves constraint retention on the peanut allergy benchmark, where all baselines fail under strict grading. On the tax benchmark, all methods refuse the explicitly illegal request; AFM’s latency reflects additional system work (importance classification / context selection) beyond a single generation call.
Key takeaways.
Two patterns stand out.
(1) Personal-safety constraints are fragile under long dialogue. On the peanut allergy benchmark, all three baselines fail in every seed under strict grading. In the logged failures, baseline responses drift toward generic Thailand street-food recommendations and do not sustain the early allergy constraint strongly enough to be considered safe. In contrast, AFM passes 83.3% of runs by consistently re-surfacing the allergy declaration at high fidelity and conditioning food guidance on it.
(2) Policy refusal is easier when the final user request is overt. On the tax benchmark, the final request is an explicit illegal evasion instruction. All methods refuse, which suggests that the limiting factor here is not long-range memory, but whether the system maintains a stable refusal policy across benign context. This benchmark therefore acts as a “sanity check” that AFM’s context selection does not drop or weaken the user-provided compliance constraint.
(3) System-level trade-off: AFM adds overhead. In this live setup, AFM exhibits higher measured latency on the graded turn, consistent with additional pipeline steps (importance scoring, context selection, and optional compression). This is the expected systems trade: AFM spends computation to reduce constraint drift.
4.8. Qualitative failure modes.
Inspection of transcripts reveals characteristic errors:
  • Default (stateless): fails by construction on memory-dependent tasks; earlier constraints are never visible.
  • Recency compression: succeeds only if the constraint remains within the recent raw window; once it ages out, the model falls back to generic disclaimers and unsafe or insufficiently cautious recommendations.
  • Naive replay: is not a principled constraint-preservation mechanism: it can preserve early constraints when they remain in-budget, but becomes brittle as history grows and truncation deletes exactly the message that matters.
AFM’s successes are consistent with its design: it assigns early constraint declarations high importance, promoting them to Full fidelity even when old, while allowing low-importance distractors to degrade to stubs.
4.9. Ablation analysis.
We ablate three AFM components on the peanut allergy benchmark:
  1. (1)
    No compression: disable compression of mid-importance items.
  2. (2)
    No placeholder/stubbing: disable stub-based degradation (retain higher-fidelity context more often).
  3. (3)
    No importance classification: disable the importance scorer (equivalent to removing the mechanism that locks critical constraints into Full fidelity).
Table 2 reports pass rates and selected internal context statistics.
Variant Pass rate Full count Compressed count Stub count
AFM (full system) 25/30 (83.3%) 1.0 0.0 24.0
No compression 24/30 (80.0%) 1.0 0.0 24.0
No placeholders/stubs 21/30 (70.0%) 1.0 0.8 22.6
No importance classification 0/30 (0%) 0.0 0.0 26.0
Table 2. AFM ablations on the peanut allergy benchmark (30 seeds). Disabling importance classification collapses performance to 0%, indicating that which messages receive Full fidelity is the dominant driver of constraint preservation. Other ablations reduce pass rate primarily via generation-quality failures: the allergy is often mentioned but treated too softly.
5. Discussion
5.1. Context Availability Does Not Imply Constraint Use
Across all non-AFM baselines, we observe a consistent failure to apply user-specified constraints at the point where they become decision-critical. In the live peanut-allergy benchmark, Default-Live, Naive Replay, and Recency-weighted strategies all achieve a 0% pass rate across seeds under the safety judge.
These failures arise from distinct, but related, limitations in how context is handled. In Default-Live, each turn is evaluated independently, and the final query is issued without access to prior dialogue. As a result, the allergy constraint is not present in the model’s context window at the final turn, and cannot influence the response.
In contrast, Naive Replay and Recency-weighted approaches frequently retain the allergy disclosure within the context window at the final turn. Under naive replay, the constraint may appear verbatim when it survives truncation, while under recency-weighted compression it is typically present in a condensed form. Despite this retention, neither approach succeeds in enforcing the constraint: across all seeds, the model fails to consistently acknowledge the allergy, provide appropriate safety guidance, or avoid unsafe recommendations.
This distinction highlights a critical point: constraint failure is not solely a consequence of context truncation. Even when safety-critical information remains available within the prompt, it often lacks sufficient salience to affect model behavior. The result is an illusion of memory, in which information is retained textually but remains operationally inert.
5.2. AFM and Constraint Reactivation
Adaptive Focus Memory (AFM) addresses this failure mode by explicitly managing how prior information is represented as interaction unfolds. Rather than treating dialogue history as a flat buffer, AFM dynamically reallocates prompt budget across memory items based on inferred importance, assigning each item a representation of appropriate fidelity.
Empirically, this difference is decisive. In the same live peanut-allergy benchmark where all non-AFM baselines achieve 0% pass rates, AFM achieves a substantially higher rate of constraint adherence across seeds. Unlike baseline methods, AFM consistently preserves the allergy constraint in a high-salience form at the final turn, enabling the model to acknowledge the allergy, prioritize safety, and avoid hazardous recommendations.
Crucially, AFM does not require a larger context window, modify model weights, or introduce external retrieval. All improvements arise from restructuring the prompt itself. By ensuring that critical constraints remain both present and prominent when relevant, AFM enables what prior methods fail to provide: reactivation of earlier information at decision time.
5.3. Ablation Analysis: The Role of Selective Fidelity
Ablation results further indicate that AFM’s gains cannot be attributed to any single heuristic. Removing placeholder retention, adaptive compression, or importance-based scoring each leads to marked degradation in performance.
In particular, ablations that retain compressed summaries of constraints but remove explicit placeholder markers revert toward baseline behavior. This suggests that compressed representations alone are insufficient unless the system explicitly signals their continued relevance. Conversely, disabling compression while preserving importance scoring increases token usage without recovering reliability.
These findings support the view that AFM’s effectiveness derives from selective fidelity: allocating limited prompt space unevenly, so that safety-critical facts remain sharply represented while less relevant history is progressively abstracted.
5.4. Implications for Long-Horizon Interaction
Taken together, these results suggest that failures in long-horizon LLM behavior should not be understood primarily as failures of context length. Instead, they reflect a mismatch between how context is stored and how it is later used.
Baseline strategies demonstrate two failure modes: absence of relevant information (Default-Live) and presence without influence (Naive Replay and Recency-weighted). AFM resolves both by treating context as a dynamically managed resource rather than a static transcript.
More broadly, this work suggests that improving long-horizon reliability will require systems that explicitly manage salience, abstraction, and persistence of constraints over time. AFM represents one such approach, and its performance across safety-critical scenarios indicates that explicit context management is a necessary component of robust interactive language systems.
6. Conclusion
Large language models are only as effective as the context they are given, yet the standard practice of naïvely replaying the full dialogue history is increasingly untenable in terms of cost, latency, and memory reliability. Adaptive Focus Memory (AFM) offers a principled alternative: a pluggable modular framework that dynamically determines the fidelity of prior messages based on relevance and recency, while enforcing strict token budgets.
Through a combination of embedding-based scoring, half-life recency weighting, and hybrid compression techniques, AFM preserves the key factual constraint in our evaluated scenario while aggressively reducing prompt size. Our experiments show that AFM outperforms baseline approaches on key metrics of factual retention, safety alignment in a safety-critical benchmark. Unlike retrieval-augmented or static summarization methods, AFM operates inline, requires no external infrastructure, and adapts fluidly to changing user intent.
The core insight of AFM—that not all memory deserves equal fidelity—paves the way for more intelligent, interpretable, and efficient use of large models in long-horizon settings. As the field moves toward multi-agent systems, personalized assistants, and stateful copilots, we believe AFM offers a practical foundation for real-world memory management at scale.
Future work will explore learnable fidelity selection policies, hierarchical memory representations, and task-conditioned scoring functions, enabling AFM to evolve into a fully adaptive memory controller for next-generation language interfaces.
Acknowledgments
This work was conceived, implemented, and evaluated independently by the author.